Dewey Corp. is expected to have an EBIT of $2,800,000 next year. Depreciation, the increase in net working capital, and capital spending are expected to be $215,000, $120,000, and $220,000, respectively. All are expected to grow at 15 percent per year for four years. The company currently has $16,500,000 in debt and 830,000 shares outstanding. After Year 5, the adjusted cash flow from assets is expected to grow at 2.7 percent indefinitely. The company’s WACC is 9.1 percent and the tax rate is 22 percent. What is the price per share of the company's stock?
Answer #1

Answer:

Expected Free Cash Flow, FCF1 = EBIT * (1 - Tax Rate) + Depreciation - Capital Spending - Investment in NWC
Expected Free Cash Flow, FCF1 = $2,800,000 * (1 - 0.22) + $215,000 - $220,000 - $120,000
Expected Free Cash Flow, FCF1 = $2,059,000

Growth rate for next 4 years is 15% and a constant growth rate (g) of 2.70% thereafter

FCF2 = $2,059,000 * 1.15 = $2,367,850
FCF3 = $2,367,850 * 1.15 = $2,723,028
FCF4 = $2,723,028 * 1.15 = $3,131,482
FCF5 = $3,131,482 * 1.15 = $3,601,204
FCF6 = $3,601,204 * 1.027 = $3,698,436

WACC = 9.10%

Horizon Value = FCF6 / (WACC - g)
Horizon Value = $3,698,436 / (0.091 - 0.027)
Horizon Value = $57,788,062.50

Value of Firm = $2,059,000/1.091 + $2,367,850/1.091^2 + $2,723,028/1.091^3 + $3,131,482/1.091^4 + $3,601,204/1.091^5 + $57,788,062.50/1.091^5
Value of Firm = $47,900,070.32

Value of Equity = Value of Firm - Value of Debt
Value of Equity = $47,900,070.32 - $16,500,000
Value of Equity = $31,400,070.32

Price per share = Value of Equity / Shares Outstanding
Price per share = $31,400,070.32 / 830,000
Price per share = $37.83